The ratio of circumradius and inradius of an equilateral triangle isA)3:1 B)1:2C)2:√3D)2:1

Grade:12

1 Answers

Yogesh Kumar
6 years ago
Let the side of the triangle be ‘a’
Circumradius=abc/4delta
incentre=delta/s
delta=area of the triangle=(\sqrt{3} x a2)/4
s=(a+a+a)/2=3a/2
Ratio: (abc/4delta)/(delta/s)=(ax 3a/2)/(4 x 3a4/16)
                                          =(3a4/2)/(3a4/4)=2/1=2:1 answer(D)